40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 13 23:39:46 2017, gautam, Update, General, ETMX suspension electronics problems? 
    Reply  Tue Jan 17 21:31:57 2017, gautam, Update, General, ETMX suspension electronics problems? 
       Reply  Wed Jan 18 10:41:14 2017, gautam, Update, General, ETMX suspension electronics problems? 
Message ID: 12716     Entry time: Fri Jan 13 23:39:46 2017     Reply to this: 12729
Author: gautam 
Type: Update 
Category: General 
Subject: ETMX suspension electronics problems? 

[Koji,gautam]

After Koji's leap second fix, we were playing around with the X arm locking. In particular, we were playing around with the limit value on the X arm LSC filter bank - the nominal value is 4000, we wanted to see if we could increase this without kicking the optic while acquiring arm lock. We initially increased it to 8000, and then turned it off altogether. Then we rapidly turned the output of the servo ON/OFF, and looked at the arm transmission to see if it came back to the level before unlocking, as an indication of whether the optic was kicked.

These trials suggested a value of 8000 for the limiter was OK, so we left the LSC mode on with the limiter set to 8000. But just as we were about to leave for the night, I noticed on the wall Striptool that the X arm was unlocked. Investigating, we found that the green wasn't even locking to a HOM. Further investigation of the Oplev spot showed that ETMX had received a large kick (both pitch and law errors were ~200urad). ITMX was unaffected.

We initially tried lowering the LSC limit value back to 4000, then used first the Oplev spot and then the green to align the arm. But turning on LSC misaligned the arm after acquiring lock. So we decided to leave LSC off, thinking that the notorious ETMX suspension problems have resurfaced. As a diagnostic, we figured we'd leave the watchdog tripped, and use the Oplev to see if the optic was getting kicked. But the act of turning the watchdog off kicked the optic again (WHY?!).

Looking at the ETMX sus screen, turning off all the damping and LSC (but watchdog on) still leaves a non-zero offset in the "Vmon" field, between 0.02-0.05V depending on the coil. Turning the watchdog OFF takes all these to 0.009V, although I can see the LR value fluctuating between 0.004V and 0.009V. I went to the Xend and squished all the cables on the Sat. Box, but the problem persisted.

At this time, I can't think of any explanation, so I am giving up for the night. To avoid unnecessarily kicking the optic, I am going to unplug the suspension from the Sat. Box and leave one of our tester boxes plugged in, lets see if that sheds any light on the situation...


Notes:

  1. The +/-20V sorensens at this end were "tripped" for a few days after the power glitch until they were reset and turned back on yesterday. But this should not affect Vmon, as these Sorensens only supply the DC voltage for the coil bias, which is a slow machine channel?
  2. The X arm was staying locked and well aligned for hours on end earlier this afternon - in fact it was locked for about 2 hours 6-8 hours ago, I can still see the trace on the wall StripTool....
ELOG V3.1.3-